# The Magnetic Filling in MAD Simulations and its Impact on the Jet in M87

### v1

@@ Abstract
Context. Magnetically arrested accretion disks (MADs) in black hole (BH) jet launching simulations are very successful in modelling low-luminosity active galactic nuclei (AGN) like M87*. The Fishbone-Moncrief (FM) torus is well established for this purpose in numerical astrophysics. The extent of the magnetic vector potential inside the FM torus - that we coin the filling factor - has not been studied before in the case of MAD simulations.
Aims. In detail, we stress the impact on the jet morphology, efficiency and linear polarization imprints of the filling factor as a significant initial parameter of these simulations.
Methods. We employ five 3D general relativistic magneto-hydrodynamics (GRMHD) simulations initialized with large-scale tori, that are immersed in weak, poloidal magnetic fields. To study the impact of the spatial extent of the initial magnetic field, hence the magnetic energy content in the torus, we scale it with the filling factor w.r.t. the poloidal geometric area of the mass density distribution. A common choice of the filling factor is complimented and investigated in terms of altered energetics and angular momentum transport. Further, we investigate the polarized, radiative imprints of synchrotron emission on M87 at 86 GHz, comparing them with Very Long Baseline Interferometry (VLBI) observations. Therefore, we perform general relativistic polarized radiative transfer calculations on the GRMHD data, modelling thermal- and non-thermal electron distributions.
Results. Our simulations show that elevated filling factors significantly increase the electromagnetic (EM) energy contributions and outward angular momentum transport in the jet, due to the initially increased magentic energy-content in the torus. High magnetic fillings exhibit increased linear polarization fractions, agreeing with the observed $\sim 15\%$ in M87*. We find the jet morphology more prone to disk-vertical flux tubes generated by MAD events. We show, that GRMHD simulations bracket the jet width measurements at the jet base in M87*.
Conclusions. Increased magnetic filling of the FM-torus produces jets that are noticeably brighter downstream compared to our reference models, hence, we find high fillings well suited for extended GRMHD jet models of other low-luminosity AGN, as well.
@@
black hole physics – magnetohydrodynamics (MHD) – accretion, accretion disks – radiative transfer – radiation mechanisms: non-thermal – globular clusters: individual: M87
@@ 1 Introduction
Messier 87 (M87) - a giant elliptical galaxy in the Virgo cluster - provides an important laboratory in the field of extra-galactic jet astrophysics as of many decades. The jet, originating from a compact radio core in the galaxy center, has been studied extensively across the electro-magnetic spectrum ranging from the radio regime up to $\gamma$-rays (e.g. Schmitt & Reid 1985; Hada et al. 2013; Kim, J.-Y. et al. 2018; Snios et al. 2019; Acciari et al. 2020; Lu et al. 2023). The EHT Collaboration et al. (EHTC) resolved the radio core of M87* at $1.3\,\mathrm{mm}$ employing the very long baseline interferometry (VLBI) technique and interpreted the ring-like structure as the gravitationally lensed synchrotron emission, originating from the accretion disk of a supermassive black hole (SMBH). In more recent observations at $86\,\mathrm{GHz}$, Lu et al. (2023) revealed that the jet and the radio core are indeed linked, which was interpreted as the jet base connecting to the accretion flow of relativistic plasma. Super-resolution images of mentioned GMVA observation and VLBA observation at $43\,\mathrm{GHz}$ were obtained, employing the resolve algortihm (cf. Kim et al. 2024a, b). The polarimetric EHT observations obtained in 2017 provide viable information on the dynamically important magnetic field structure. From these observations a linear polarization fraction of $\sim 15\,\mathrm{\%}$ EHT Collaboration et al. (2021a) and an upper bound on the circularly polarized emission fraction of $\langle|\nu|\rangle<3.7\%$ EHT Collaboration et al. (2023) was found. These results constrain theoretical models such as general relativistic magneto-hydrodynamics (GRMHD) simulations of the plasma flow in the vicinity of the black hole. These simulations can be compared to observations by means of general relativistic radiative transfer (GRRT) post-processing calculations on the GRMHD data. Yang et al. (2024) found the 86 GHz and 43 GHz jet morphology of similar GRMHD simulations to fit the observations very well. Cruz-Osorio et al. (2022) and Fromm et al. (2022) found high spin parameters of $a\sim 0.5-0.94$ to be favourable to match the broad-band spectrum and jet morphology of M87* in wide parameter space studies of GRMHD models. In this work we want to focus on the comparison of five 3D GRMHD models and their impact on the electromagnetic footprint of the M87 jet. All share similar initial conditions, but differ with respect to their initial magnetic field configuration. We use the Fishbone-Moncrief (FM) torus Fishbone & Moncrief (1976) and initialize it with a weak poloidal magnetic field. In most studies, the poloidal extent of the initial magnetic field is set with the offset of the initial vector-potential from a zero-cutoff to $0.80$. We compare this model with four simulations that fill the initial torus with varying spatial extents of a weak poloidal magentic field. These models exhibit drastically altered energetic flux, angular momentum transport properties and jet generation efficiencies. Chatterjee & Narayan (2022) studied the behavior of and drivers for angular momentum transport in magnetically arrested disk (MAD) simulations compared to standard and normal evolution (SANE) simulations in non-spinning BH systems to isolate the side-effects of frame dragging. We compliment this study with the case of BHs with high spin and compare the impact of the magnetic filling of the torus. Recent studies by Jacquemin-Ide et al. (2025) complement explanations for the underlying mechanism of MAD-states by studying the inward advective and outward diffusive cycles these simulations exhibit and based on that, derive a first analytical expression for the recurrence timescale of MAD-states. Further parametrical investigation of the MAD-setup by Cho & Narayan (2025) found that the overall variability in the simulations is sensitive to the ratio of rotational and magnetic energy in the initial setup, which connects well to our study, where we tune the initial magnetic energy content.
@@
This paper contains two major parts: In Sec. 2 we introduce the underlying GRMHD equations and simulation setups. We conduct tests on the resolution of the magneto-rotational instability (MRI). Thereafter, we study the differences in the jet energetics and will then determine the efficiency of the energy extraction from the BH and disk to the energy flux in the jet. We connect this analysis further with the investigation of the angular momentum transport of the five simulations. Connecting to the morphological study of the radiative propertities of a subset of three simulations in Sec. 3, we consider the impact of magnetic flux eruptions from the BH on the hydromdynamical jet structure, of these simulations. In Sec. 3 we ray-trace the polarized synchrotron emission from a population of thermal and non-thermal electrons asserted to the GRMHD simulations. After the ray-tracing setup is introduced, we will compare the jet morphology at 86 GHz and present spatially resolved, as well as time-dependent linear polarization fractions. Next we compare the jet edge and width profiles of the three different models with observations of Lu et al. (2023) and discuss the impact of flux eruptions on the synchrotron emission from the jet and disk.
2 General Relativistic Magneto-Hydrodynamics
@@
We employ the 3D GRMHD code Kokkos-based High-Accuracy Relativistic Magnetohydrodynamics with AMR (KHARMA; Prather et al. (2021); Prather (2024)), that simulates accretion onto and jet launching from black holes. KHARMA solves the ideal, non-radiative MHD equations on curved space-times with four-metric $g_{\mu\nu}$ and metric determinant $g$ in geometric units ($G=c=1$ and a factor of $\nicefrac{{1}}{{\sqrt{4\pi}}}$ is absorbed in the magnetic field). Throughout this paper we will use the code units for the time- and length-scales, i.e. the light-crossing time is redifend to $t_{\mathrm{g}}=1\,G_{\mathrm{N}}M_{\mathrm{BH}}/(c^{3})\equiv 1\,\mathrm{M}$ and the gravitational radius is $r_{\mathrm{g}}=1\,G_{\mathrm{N}}M_{\mathrm{BH}}/(c^{2})\equiv 1\,\mathrm{M}$, where $G_{\mathrm{N}}$ is Newton’s constant, $c$ is the vacuum speed of light, $M_{\mathrm{BH}}$ is the BH mass and $\mathrm{M}$ is the mass unit that scales the units in the simulations. Greek indices run through $(0,1,2,3)$ and Roman indices through $(1,2,3)$ and we make use the Einstein summation convention. We use a Kerr black hole solution on a modified spherical coordinate system (Funky-Modified-Kerr-Schild-coordinates; FMKS). These coordinates exhibit a decreasing grid cell size in $\theta$-direction towards the equator ($\theta=\nicefrac{{\pi}}{{2}}$), that become increasingly cylindrical along the jet-axis ($\theta=0$ and $\theta=\pi$). The radial extent of the cells increase exponentially, therefore high resolution is given at small radii, where finely resolving the accretion process is vital. ($r$, $\theta$, $\phi$) refer to the ordinary spherical coordinates. The simulation grids are equal in all simulations, where we employ $N_{r}=256$, $N_{\phi}=128$ and $N_{\theta}=128$ cells in the respective directions. The radial boundaries range from $r=(1.1659-1000)\,\mathrm{M}$, and otherwise use the complete map in $\phi=(0-2\pi)\,\mathrm{rad}$ and $\theta=(0-\pi)\,\mathrm{rad}$. The underlying equations solved by KHARMA are the covariant conservation of mass, local conservation of energy-momentum and the covariant Maxwell equations (cf. Mizuno & Rezzolla (2024) for more details on the derivation):
@@
with total specific enthalpy $h_{\mathrm{tot}}=h+\nicefrac{{b^{2}}}{{\rho}}$ and fluid pressure $p$. The magnetic field four-vector $b^{\mu}$ in its temporal component and spatial components reads:
@@
where $B^{i}$ is the magnetic field three-vector measured by an Eulerian observer, $\alpha$ the lapse function and $\Gamma$ the bulk Lorentz factor. Note, that with the definition of $b^{\mu}$ in Eq. (3), we make use of the notation $b^{2}=b^{\mu}b_{\mu}$ in Eq. (2). We use the plasma-beta $\beta=\nicefrac{{p}}{{b^{2}}}$ and magnetization $\sigma=\nicefrac{{b^{2}}}{{\rho}}$ parameters for diagnostics with:
$b^{2}=\frac{B^{2}+\alpha^{2}(b^{t})^{2}}{\Gamma^{2}}=\frac{B^{2}}{\Gamma^{2}}+(B^{i}v_{i})^{2}$ (4)
The system of equations is closed by an equation of state. We assume an ideal gas, that relates the enthalpy $h$ to the pressure $p$ and density $\rho$:
@@
with the adiabatic index $\hat{\gamma}$ (Antón et al. (2006)).
We employ inflow and outflow boundary conditions in a radial coordinate, along the polar boundaries we use a solid reflective wall (Olivares et al. (2019)). In the azimuthal direction we employ periodic boundary conditions to all physical quantities across the cells at $\phi=0$. The GRMHD equations are solved with finite-volume and high-resolution, shock-capturing methods, that are the Lax-van Leer Friedrichs (LLF) flux formula in combination with the WENO5 reconstruction scheme and in synchronization with the IMEX scheme. The density floor is set to $\rho_{\mathrm{min}}=10^{-6}$, the internal energy floor to $u_{\mathrm{min,geo}}=10^{-8}$, the ceiling of the magnetization is $\sigma_{\mathrm{max}}=100$ and the ceiling of the internal energy is $\left(\nicefrac{{u}}{{\rho}}\right)_{\mathrm{max}}=2$. The initial pressure maximum in the torus is set to $\rho_{\rm{max}}=1$. The 3D GRMHD simulations are initialized with a magnetized torus in hydrodynamic equilibrium with a rotating Kerr BH in the center of the simulation domain. Here, we use a setup that allows the disk to partially become magnetically arrested. In the MAD setup we add a single poloidal loop in the vector potential on top of the FM torus (Fishbone & Moncrief (1976)) in the following way:
@@
$q=\frac{\rho}{\rho_{\mathrm{max}}}\left(\frac{r}{r_{\mathrm{in}}}\right)^{3}\sin^{3}\hskip-2.0pt\theta\,\exp{\left(\frac{-r}{400}\right)},$ (7)
@@
Figure 1: Invariant magnetic field strength $b$, the spatial magnetic field components $B^{r}$ and $B^{\theta}$, and four-vector-potential component $A_{\phi}$ over the polar angle at the initialization step ($t=0\,\rm{M}$). Colors correspond to simulations K.60 (orange), K.80 (green), K.90 (blue), K.95 (magenta) and K.99 (red). In the middle row the solid lines show $B^{r}$ and the dashed lines $B^{\theta}$. We choose polar slices at radii around the inner edge of the torus at $r_{\rm{in}}=20\,\rm{M}$ and at the pressure maximum in the tori at $r=50\,\rm{M}$. Note that going from left to right colums, i.e. radially outwards, the simulations with higher filling factors show increased magnetic field strengths and extended magnetic field, both radially and in the polar-angle, especially at the inner edge of the torus.
@@
We solely vary the filling factor $F$ throughout the simulations from 0.60.990.6-0.99, keeping other initial conditions of the GRMHD simulations constant. Those are constant adiabatic index $\hat{\gamma}=\frac{5}{3}$, inner radius $r_{\mathrm{in}}=20$ and radius of the density maximum $r_{\mathrm{c}}=41$ of the torus as well as the dimensionless spin parameter $a=15/16$ of the BH. The plasma-beta is initialized such that its minimal value is above $\beta_{\rm{min}}=2\,\nicefrac{{p}}{{b^{2}}}=100$. We refer to the five models as K.60, K.80, K.90, K.95 and K.99 corresponding to the simulations in that the filling factor is set to $F=0.60$, $F=0.80$, $F=0.90$, $F=0.95$ and $F=0.99$, respectively. In Appendix A we present more details on the correspondance of the filling factor to the volumes initialized with magnetic field relative to the torus volume.
@@
In Eq. (6), the effect of $F$ is to proportionally scale the off-set from the zero cut-off of $A_{\phi}$, that is visible in the bottom row of Fig. 1. With increasing $F$, the off-set of $A_{\phi}$ increases, which in turn increases the poloidal extent, or to be precise, the total volume containing cells that have non-zero magnetic field values. For further details we refer to Appendix A, where we also address that the filling factor does not scale the poloidal-field extent linearly to the extent of the torus, it is rather close to the latter and for $F\gtrsim 0.80$ above it (cf. Appendix Tab. 3).
@@
Figure 2: Mass accretion rates $\dot{M}$, magnetic flux through the event horizon of the BH $\phi_{\mathrm{BH}}$ and the MAD-parameter $\Psi=\phi_{\mathrm{BH}}/\sqrt{\dot{M}}$ for the simulations K.80, K.90 and K.99 over time.
@@
All simulations are calculated up to a simulation time of $30000\,\mathrm{M}$. We show the mass accretion rate $\dot{M}$, the magnetic flux through the event horizon $\phi_{\mathrm{BH}}$ and the MAD-parameter $\Psi=\nicefrac{{\phi_{\mathrm{BH}}}}{{\sqrt{\dot{M}}}}$ over the full temporal range of the simulations in Fig. 2.
In Appendix B Fig. 16 we present azimuthal and time-averaged cross-sections of the meridional plane of $\rho$, $\sigma$, magnetic field magnitude $b=\sqrt{|b_{\mu}b^{\mu}|}$, plasma-$\beta$ and ion temperature $\Theta_{i}$.

### v2

# The magnetic filling in magnetically arrested accretion disk simulations and its impact on the jet in M87

@@ Abstract
Context. Magnetically arrested accretion disks (MADs) in black-hole (BH) jet-launching simulations are very successful in modeling low-luminosity active galactic nuclei (AGNs) such as M87*. The Fishbone-Moncrief (FM) torus is well established for this purpose in numerical astrophysics. The extent of the magnetic vector potential inside the FM torus –which we dub the filling factor– has not been studied before in the case of MAD simulations.
Aims. In detail, we stress the impact on the jet morphology, efficiency, and linear polarization imprints of the filling factor as a significant initial parameter of these simulations.
Methods. We employed five 3D, general-relativistic magnetohydrodynamic (GRMHD) simulations initialized with large-scale tori, which are immersed in weak, poloidal magnetic fields. To study the impact of the spatial extent of the initial magnetic field, and hence the magnetic energy content in the torus, we scaled it with the filling factor with regard to the poloidal geometric area of the mass-density distribution. A common choice of the filling factor is complimented and investigated in terms of altered energetics and angular-momentum transport. Further, we investigated the polarized, radiative imprints of synchrotron emission on M87 at 86 GHz, comparing them with very long baseline interferometry (VLBI) observations. Therefore, we performed general-relativistic, polarized radiative transfer calculations on the GRMHD data, modeling thermal and nonthermal electron distributions.
Results. Our simulations show that elevated filling factors significantly increase the electromagnetic (EM) energy contributions and outward angular-momentum transport in the jet due to the initially increased magnetic energy content in the torus. High magnetic fillings exhibit increased linear polarization fractions, agreeing with the observed $\sim 15\%$ in M87*. We find the jet morphology more prone to disk vertical flux tubes generated by MAD events. We show that GRMHD simulations bracket the jet width measurements at the jet base in M87*.
Conclusions. Increased magnetic filling of the FM torus produces jets that are noticeably brighter downstream compared to our reference models; hence, we also find high fillings well suited for extended GRMHD jet models of other low-luminosity AGNs.
@@
black-hole physics – magnetohydrodynamics – accretion, accretion disks – radiative transfer – radiation mechanisms: nonthermal – globular clusters: individual: M87
@@ 1 Introduction
Messier 87 (M87), a giant elliptical galaxy in the Virgo cluster, has provided an important laboratory in the field of extragalactic jet astrophysics for many decades. The jet, originating from a compact radio core in the galaxy center, where the active galactic nucleus (AGN) is located, has been studied extensively across the electromagnetic spectrum, ranging from the radio regime up to $\gamma$-rays (e.g. Schmitt and Reid 1985; Hada et al. 2013; Kim, J.-Y. et al. 2018; Snios et al. 2019; Acciari et al. 2020; Lu et al. 2023). The Event Horizon Telescope Collaboration (EHTC; EHTC et al. (2019a)) resolved the radio core of M87* at $1.3\,\mathrm{mm}$ employing the very long baseline interferometry (VLBI) technique and interpreted the ring-like structure as the gravitationally lensed synchrotron emission, originating from the accretion disk of a supermassive black hole (SMBH). In more recent observations at $86\,\mathrm{GHz}$, Lu et al. (2023) revealed that the jet and the radio core are indeed linked, which was interpreted as the jet base connecting to the accretion flow of relativistic plasma. Super-resolution images of mentioned observations with the Global Millimetre VLBI Array (GMVA) and Very Long Baseline Array (VLBA) at $43\,\mathrm{GHz}$ were obtained, employing the resolve algorithm (see Kim et al. 2025, 2024). The polarimetric EHT observations obtained in 2017 provide viable information on the dynamically important magnetic-field structure. From these observations, a linear polarization fraction of $\sim 15\,\mathrm{\%}$ EHTC et al. (2021a) and an upper bound on the circularly polarized emission fraction of $\langle|\nu|\rangle<3.7\%$ EHTC et al. (2023) was found. These results constrain theoretical models such as general-relativistic magnetohydrodynamic (GRMHD) simulations of the plasma flow in the vicinity of the black hole (BH). These simulations can be compared to observations by means of general-relativistic radiative-transfer (GRRT) post-processing calculations on the GRMHD data. Yang et al. (2024) found the 86 GHz and 43 GHz jet morphology of similar GRMHD simulations to fit the observations very well. Cruz-Osorio et al. (2022) and Fromm et al. (2022) found high spin parameters of $a\sim 0.5-0.94$ to be favorable to match the broad-band spectrum and jet morphology of M87* in wide parameter-space studies of GRMHD models. In this work, we focused on the comparison of five 3D GRMHD models and their impact on the electromagnetic footprint of the M87 jet. All share similar initial conditions, but differ with respect to their initial magnetic-field configuration. We used the Fishbone-Moncrief (FM) torus Fishbone and Moncrief (1976) and initialized it with a weak poloidal magnetic field. In most studies, the poloidal extent of the initial magnetic field is set with the offset of the initial vector potential from a zero cutoff to $0.80$. We compared this model with four simulations that fill the initial torus with varying spatial extents of a weak poloidal magnetic field. These models exhibit drastically altered energetic-flux, angular-momentum transport properties and jet-generation efficiencies. Chatterjee and Narayan (2022) studied the behavior and drivers of angular-momentum transport in magnetically arrested disk (MAD) simulations compared to standard- and normal-evolution (SANE) simulations in non-spinning BH systems to isolate the side effects of frame dragging. We complemented this study with the case of BHs with high spin and compared the impact of the magnetic filling of the torus. Recent studies by Jacquemin-Ide et al. (2025) complement explanations for the underlying mechanism of MAD states by studying the inward advective and outward diffusive cycles these simulations exhibit, and, based on that, they derived a first analytical expression for the recurrence timescale of MAD states. Further parametrical investigation of the MAD setup by Cho and Narayan (2025) found that the overall variability in the simulations is sensitive to the ratio of rotational and magnetic energy in the initial setup, which connects well to our study, where we tuned the initial magnetic-energy content.
@@
This paper contains two major parts. In Sect. 2, we introduce the underlying GRMHD equations and simulation setups. We describe the tests we conducted on the resolution of the magneto-rotational instability (MRI). Thereafter, we explain our study of the impact of the initial magnetic energy in the torus on the jet energetics and on the efficiency of energy extraction from the BH and disk to the energy flux in the jet. We connect this analysis further with the investigation of the angular momentum transport of the five simulations. In connecting it to the morphological study of the radiative propertities of a subset of three simulations in Sect. 3, we considered the impact of magnetic-flux eruptions from the BH on the hydrodynamical jet structure of these simulations. In Sect. 3, we describe how we ray-traced the polarized synchrotron emission from a population of thermal and nonthermal electrons asserted by the GRMHD simulations. After the ray-tracing setup is introduced, we compare the jet morphology at 86 GHz and present spatially resolved, as well as time-dependent, linear polarization fractions. Finally, we compare the jet edge and width profiles of the three different models with observations from Lu et al. (2023) and discuss the impact of flux eruptions on the synchrotron emission from the jet and disk.
2 General-relativistic magnetohydrodynamics
@@
We employed the 3D GRMHD code Kokkos-based High-Accuracy Relativistic Magnetohydrodynamics with AMR (KHARMA; Prather et al. (2021); Prather (2024)), which simulates accretion onto and jet launching from BHs. KHARMA solves the ideal magneto-hydrodynamics (MHD) equations on curved space times with a four-metric $g_{\mu\nu}$ and a metric-determining $g$ in geometric units ($G=c=1,$ and a factor of $\nicefrac{{1}}{{\sqrt{4\pi}}}$ is absorbed in the magnetic field). Throughout this paper, we use code units for the time and length scales; i.e., the light-crossing time is redefined as $t_{\mathrm{g}}=1\,G_{\mathrm{N}}M_{\mathrm{BH}}/(c^{3})\equiv 1\,\mathrm{M,}$ and the gravitational radius is $r_{\mathrm{g}}=1\,G_{\mathrm{N}}M_{\mathrm{BH}}/(c^{2})\equiv 1\,\mathrm{M}$, where $G_{\mathrm{N}}$ is Newton’s constant, $c$ is the vacuum speed of light, $M_{\mathrm{BH}}$ is the BH mass, and $\mathrm{M}$ is the mass unit that scales the units in the simulations. Greek indices run through $(0,1,2,3)$ and Roman indices through $(1,2,3),$ and we made use of the Einstein summation convention. We used a Kerr black-hole solution on a modified spherical coordinate system including funky-modified–Kerr–Schild coordinates (FMKS). These coordinates exhibit a decreasing grid cell size in the $\theta$-direction toward the equator ($\theta=\nicefrac{{\pi}}{{2}}$) that becomes increasingly cylindrical along the jet-axis ($\theta=0$ and $\theta=\pi$). The radial extent of the cells increases exponentially; therefore, high resolution is given at small radii, where finely resolving the accretion process is vital. ($r$, $\theta$, $\phi$) refer to the ordinary spherical coordinates. The simulation grids are equal in all simulations, in which we employed $N_{r}=256$, $N_{\phi}=128,$ and $N_{\theta}=128$ cells in the respective directions. The radial boundaries range from $r=(1.1659-1000)\,\mathrm{M}$ and otherwise use the complete map in $\phi=(0-2\pi)\,\mathrm{rad}$ and $\theta=(0-\pi)\,\mathrm{rad}$. The underlying equations solved by KHARMA are the covariant conservation of mass, local conservation of energy-momentum, and the covariant Maxwell equations (see Mizuno and Rezzolla (2024) for more details on the derivation):
@@
with a total specific enthalpy of $h_{\mathrm{tot}}=h+\nicefrac{{b^{2}}}{{\rho}}$ and a fluid pressure of $p$. The magnetic-field four-vector $b^{\mu}$ in its temporal component and spatial components reads
@@
where $B^{i}$ are the spatial components of the magnetic-field three-vector measured by an Eulerian observer, $\alpha$ is the lapse function, and $\Gamma$ is the bulk Lorentz factor. Note that with the definition of $b^{\mu}$ in Eq. (3), we made use of the notation $b^{2}=b^{\mu}b_{\mu}$ in Eq. (2). We used the plasma beta, $\beta=\nicefrac{{2p}}{{b^{2}}}$, and magnetization, $\sigma=\nicefrac{{b^{2}}}{{\rho}}$, parameters for diagnostics with
$b^{2}=\frac{B^{2}+\alpha^{2}(b^{t})^{2}}{\Gamma^{2}}=\frac{B^{2}}{\Gamma^{2}}+(B^{i}v_{i})^{2}.$ (4)
The system of equations is closed by an equation of state. We assumed an ideal gas that relates the enthalpy, $h,$ to the pressure, $p,$ and density, $\rho$:
@@
where the adiabatic index is $\hat{\gamma}$ (Antón et al. (2006)).
We employed inflow and outflow boundary conditions in a radial coordinate. Along the polar boundaries, we used a solid reflective wall (Olivares et al. (2019)). In the azimuthal direction, we employed periodic boundary conditions to all physical quantities across the cells at $\phi=0$. The GRMHD equations were solved with finite-volume and high-resolution shock-capturing methods including the Lax-van Leer Friedrichs (LLF) flux formula in combination with the WENO5 reconstruction scheme and in synchronization with the IMEX scheme. The density floor was set to $\rho_{\mathrm{min}}=10^{-6}$ and the internal energy floor to $u_{\mathrm{min,geo}}=10^{-8}$; the ceiling of the magnetization was $\sigma_{\mathrm{max}}=100,$ and the ceiling of the internal energy was $\left(\nicefrac{{u}}{{\rho}}\right)_{\mathrm{max}}=2$. The initial pressure maximum in the torus was set to $\rho_{\rm{max}}=1$. The 3D GRMHD simulations were initialized with a magnetized torus in hydrodynamic equilibrium with a rotating Kerr BH in the center of the simulation domain. Here, we used a setup that allows the disk to become partially magnetically arrested. In the MAD setup, we added a single poloidal loop in the vector potential on top of the FM torus (Fishbone and Moncrief (1976)) in the following way:
@@
$q=\frac{\rho}{\rho_{\mathrm{max}}}\left(\frac{r}{r_{\mathrm{in}}}\right)^{3}\sin^{3}\hskip-2.0pt\theta\,\exp{\left(\frac{-r}{400}\right)}.$ (7)
@@
Figure 1: Invariant magnetic-field strength, $b$; the spatial magnetic-field components, $B^{r}$ and $B^{\theta}$; and four-vector-potential component, $A_{\phi},$ over the polar angle at the initialization step ($t=0\,\rm{M}$). Colors correspond to simulations K.60 (orange), K.80 (green), K.90 (blue), K.95 (magenta), and K.99 (red). In the middle row, the solid lines show $B^{r}$ and the dashed lines $B^{\theta}$. We chose polar slices at radii close to the inner edge of the torus at $r_{\rm{in}}=20\,\rm{M}$ and at the pressure maximum in the tori at $r=50\,\rm{M}$. Note that in columns going from left to right, i.e., radially outwards, the simulations with higher filling factors show increased magnetic-field strengths and extended magnetic field, both radially and in the polar-angle, especially at the inner edge of the torus.
@@
We solely varied the filling factor, $F,$ throughout the simulations from 0.60.990.6-0.99, keeping other initial conditions of the GRMHD simulations constant. Those are a constant adiabatic index, $\hat{\gamma}=\frac{5}{3}$; inner radius, $r_{\mathrm{in}}=20;$ and radius of the density maximum, $r_{\mathrm{c}}=41,$ of the torus as well as the dimensionless spin parameter, $a=15/16,$ of the BH. The plasma beta was initialized such that its minimal value was above $\beta_{\rm{min}}=2\,\nicefrac{{p}}{{b^{2}}}=100$. We refer to the five models as K.60, K.80, K.90, K.95, and K.99; these correspond to the simulations in that the filling factor is set to $F=0.60$, $F=0.80$, $F=0.90$, $F=0.95,$ and $F=0.99$, respectively. In Appendix A, we present more details on the correspondance of the filling factor to the volumes initialized with the magnetic field relative to the torus volume.
@@
In Eq. (6), the effect of $F$ is to proportionally scale the offset from the zero cutoff of $A_{\phi}$, which is visible in the bottom row of Fig. 1. With increasing $F$, the offset of $A_{\phi}$ increases, which in turn increases the poloidal extent or, to be precise, the total volume containing cells that have nonzero magnetic-field values. For further details, we refer the reader to Appendix A, where we also address that the filling factor does not scale the poloidal-field extent linearly to the extent of the torus; it is rather close to the latter, and for $F\gtrsim 0.80$ it is above it (see Tab. 3).
@@
Figure 2: Mass-accretion rates, $\dot{M}$, and magnetic flux through the event horizon of the BH, $\phi_{\mathrm{BH,}}$ and the MAD-parameter, $\Psi=\phi_{\mathrm{BH}}/\sqrt{\dot{M,}}$ for all simulations over time.
@@
All simulations were calculated up to a simulation time of $30000\,\mathrm{M}$. We show the mass accretion rate, $\dot{M}$; the magnetic flux through the event horizon, $\phi_{\mathrm{BH}}$; and the MAD parameter, $\Psi=\nicefrac{{\phi_{\mathrm{BH}}}}{{\sqrt{\dot{M}}}}$, over the full temporal range of the simulations in Fig. 2. In Appendix B, Fig. 16, we present azimuthal and time-averaged cross-sections of the meridional plane of $\rho$, $\sigma$, magnetic field magnitude $b=\sqrt{|b_{\mu}b^{\mu}|}$, plasma-$\beta,$ and the ion temperature, $\Theta_{i}$.